# Enhancing Negation Awareness in Universal Text Embeddings: A Data-efficient and Computational-efficient Approach

### v1

@@ Abstract
Negation plays an important role in various natural language processing tasks such as Natural Language Inference and Sentiment Analysis tasks. Numerous prior studies have found that contextual text embedding models such as BERT, ELMO, RoBERTa or XLNet face challenges in accurately understanding negation. Recent advancements in universal text embeddings have demonstrated superior performance over contextual text embeddings in various tasks. However, due to the bias in popular evaluation benchmarks, the negation awareness capacity of these models remains unclear. To bridge the gap in existing literature, an in-depth analysis is initiated in this work to study the negation awareness of cutting-edge universal text embedding models. Our findings reveal a significant lack of negation awareness in these models, often interpreting negated text pairs as semantically similar. To efficiently deal with the conflict that different tasks need different trade-offs between topic and negation information among other semantic information, a data-efficient and computational-efficient embedding re-weighting method is proposed without modifying the parameters of text embedding models. The proposed solution is able to improve text embedding models’ negation awareness significantly on both simple negation understanding task and complex negation understanding task. Furthermore, the proposed solution can also significantly improve the negation awareness of Large Language Model based task-specific high dimensional universal text embeddings.
Index Terms:
Negation awareness, Semantic similarity, Universal text embeddings, Natural Language Processing
I Introduction
@@
Text embedding has gained significant attention from both industry and academia due to its crucial role in various Natural Language Processing (NLP) tasks such as information retrieval [1], sentiment analysis [2, 3], text classification [4], text clustering [5, 6], etc. Recently, the importance of text embeddings has been further highlighted in the context of Large Language Models (LLMs) based applications, such as Retrieval Augmented Generation (RAG) systems. This is primarily because these applications rely on high-quality text embeddings for performing vector search, which involves retrieving the most relevant contexts/documents for LLM Question Answering (QA) tasks [7].
One crucial aspect related to the effectiveness of text embedding model is the comprehension and interpretation of negation information, a key linguistic construct that significantly influences sentiment and meaning [8]. Its accurate interpretation is therefore critical in sentiment analysis and Natural Language Inference (NLI) tasks [9, 10]. However, several previous works have shown that contextual text embeddings like BERT [11] lack understanding of negations and fail to attribute sufficient importance to the word “not” [12, 13, 10]. Such bias could render the entire output of the model invalid, thereby diminishing the reliability of such systems, particularly in scenarios involving opinion mining, fact-checking or measuring answer correctness [8].
@@
The field of universal text embeddings has recently witnessed remarkable advancements. This progress can be attributed to the advancements in the quantity and quality of diverse text datasets across different tasks [14, 15], synthetic data generated by different LLMs [16, 17], using LLMs as backbones and the emergence of benchmarks with the focus on novel task and domain generalization such as the Massive Text Embedding Benchmark (MTEB) [18]. Representative universal embedding models such as GTE [19], BGE [14], E5 [20, 17, 21], Gecko [16] or LLM2Vec [22] can effectively address diverse downstream tasks without fine-tuning. They have demonstrated significant improvements over contextual text embeddings across diverse tasks including information retrieval, reranking, clustering and pair classification tasks [7]. However, evaluation benchmarks such as MTEB are biased towards tasks such as retrieval, classification and clustering [18, 7, 23]. The negation awareness of state-of-the-art universal text embeddings remains unclear.
@@
  • Novel in-depth analysis is proposed to understand how similar are negated text pairs interpreted by diverse state-of-the-art universal text embeddings.
  • To deal with the conflict that different tasks need different trade-offs between topic and negation information among other semantic information, a data-efficient and sustainable embedding re-weighting method is proposed without fine-tuning.
@@
II Related Works
II-A Universal text embeddings
The field of text embeddings has seen substantial improvements in recent years. Text embedding models can be categorized in four eras [7]: the first era stands for count-based embeddings such as Bag of Words [24] and Term Frequency-Inverse Document Frequency (TF-IDF) [25] which are transformed into low-dimensional dense embeddings with methods like Latent Semantic Indexing (LSA) [26]; the second era is static dense word embeddings represented by Word2Vec [27], GloVe [28] and FastText [29] which can not deal with polysemy; the third era is contextualized embeddings where context-sensitive dynamic embeddings can adapt or change based on context and deal with polysemy, represented by Embeddings From Language Models (ELMo) [30], Generative Pre-trained Transformer (GPT) [31] and Bidirectional Encoder Representations from Transformers (BERT) [11]. Recently, there has been significant progress in the fourth generation of universal text embeddings which aim at building an unified comprehensive text embedding to address a multitude of input text length, downstream tasks, domains and languages [7].
@@ II-B Negation understanding
Negation plays an important role in many NLP tasks including NLI [8], sentiment analysis [9], Natural Language Understanding (NLU) tasks [10], etc. The complexity of negation, as highlighted by [9], presents substantial challenges for commercial sentiment analysis. Contextual text embedding models such as BERT [11] is shown to be insensitive to negations in [12] through a completion task involving masked hypernyms in affirmative and negated versions of each sentence. Other Contextual text embeddings including Transformer-XL [36] and ELMO [30] are also found to exhibit poor performance in distinguishing between affirmative and negative sentences when subjected to negative cloze tests [37]. Moreover, they tend to make identical predictions regardless of the polarity of the sentence [38]. In the study conducted by [13, 10], novel benchmark was created to evaluate the ability of contextual embedding models in terms of recognizing negation. The experimental findings revealed that even after fine-tuning with additional negation-containing pairs, contextual text embeddings (RoBERTa [39], XLNet [40] and BERT [11]) still face challenges in accurately understanding negation.
Several studies have been conducted in order to improve the negation awareness of contextual text embedding models such as BERT. In [41], BERTNOT is proposed by fine-tuning a pre-trained BERT model using a knowledge distillation objective as well as a novel unlikelihood objective, which is based on negated generic sentences extracted from a raw text corpus. The proposed BERTNOT has been observed to effectively incorporate negation knowledge into the BERT-base model. However, it has been found to be ineffective with the BERT-large model. A novel negation-focused pre-training strategy is proposed in [42] to improve the performance and generalization ability of negation detection using targeted data augmentation and negation masking. Another new pre-training framework with a weighted cross-entropy loss and elastic weight consolidation regularization was proposed in [43] to overcome catastrophic forgetting and augment the negation knowledge in contextual text embedding models.
@@
III How similar are negated text pairs for universal text embeddings?
@@
Semantic similarity refers to the degree of overlap in meaning [35]. Measuring the semantic similarity using cosine over text embeddings is one of the most adopted method. Among the 8 evaluation tasks from MTEB [18], embedding based semantic similarity is used by 7 of them. Recently, text embedding based semantic similarity is also used as evaluation metrics to measure the answer correctness of RAGs or LLMs [44, 45]. Despite its wide usage, there is few study trying to understand what kind of semantic information is encoded and highlighted in the semantic similarity based on text embeddings. In the literature of Semantic Textual Similarity (STS) studies such as [46, 35], semantic similarity is defined as (re-scaled to the range between 0 and 1):
  • 0 means the the pair of texts are on different topics;
  • 0.2 means the the pair of texts are not equivalent, but are on the same topic;
  • 0.4 means the the pair of texts are not equivalent, but share some details;
@@
  • 0.6 means the the pair of texts are roughly equivalent, but some important information differs/missing;
@@
  • 0.8 means the the pair of texts are mostly equivalent, but some unimportant details differ;
  • 1 means the the pair of texts are completely equivalent;
@@
One limitation of this definition is that: the semantic similarity is mainly based on topic information while other semantic information such as negation is not included. When it comes to two pieces of text where one is the negated version of the other, the situation becomes more complex. For instance, consider these two sentences: (1) ”The horse is white.” and (2) ”The horse is not white.”: from a purely syntactic perspective, the sentences are very similar. They share the same subject and predicate, with the only difference being the presence of negation in (2). From a purely topic perspective, the sentences are similar as they share similar topics: the color of the horse. However, semantic textual similarity is not just about the structure or the words used, but about the meaning conveyed by the sentences. The negation in the second sentence changes the meaning entirely. According to [43], the significance of negation in understanding the semantic relationship in natural language texts is profound, as it reverses the overall semantic meaning of a sentence. Consequently, the primary focus of this section is to study how universal text embeddings interpret the semantic similarity between negated text pairs.
III-A Dataset
@@
STSB [35] is a collection of English datasets that have been used in the SemEval and *SEM STS shared tasks over the period from 2012 to 2017 [35]. Annotation of text pair’s similarity is accomplished via crowdsourcing, integrating pragmatic and global knowledge, thereby enhancing their interpretability and utility for subsequent tasks [35]. Within the scope of this study, the human annotation similarity scores in the STSB dataset, which typically range from 0 to 5, are re-scaled to the range between 0 and 1. Given STSB text pairs with the human annotated semantic similarity in the form of [$sentence\_1_{i}$, $sentence\_2_{i}$, $similarity\_score_{i}$], a sentence-level negation tool (focuses on verbal negations and supports the simple addition and deletion of negation cues) developed by [8] is used to create $neg\_sentence\_1_{i}$: the simple negated version of $sentence\_1_{i}$ (e.g. transform ”The horse is white” into its simple negated version of ”The horse is not white”). The training partition of the dataset is used for analysis (in this section) and training (in next section), while the development and test sets are combined for testing (in next section).
III-B Universal text embeddings
@@ III-C Experiments
In order to study how universal text embeddings interpret negation, the training data of STSB are separated into five groups based on the human annotated similarity between $sentence\_1_{i}$ and $sentence\_2_{i}$: group 1 (0-0.2), group 2 (0.2-0.4), group 3 (0.4-0.6), group 4 (0.6-0.8), and group 5 (0.8-1). The semantic similarity $Sim\_12$ between S1 ($sentence\_1$) and S2 ($sentence\_2$) as well as the semantic similarity $Sim\_1neg1$ between S1 and its negated version NegS1 ($neg\_sentence\_1$) within each group based on each universal text embeddings are calculated. By comparing $Sim\_1neg1$ with $Sim\_12$ relatively within each group, we can get the insights on how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings.
@@
The histogram of $Sim\_12$ within each group is illustrated in blue color in Figure 2: row 1 to row 5 represents the results from group 1 to group 5 and each column represents a text embedding model. Generally speaking, the absolute similarity values do not match the human annotation. For instance, the human annotation for group 1 is between 0 and 0.2, while the similarity values calculated by most embedding models are between 0.2 and 0.8. For E5 series, similarity values are between 0.6 and 1 for group 1. However, the evaluation metric used by STSB is Pearson correlation between embedding based semantic similarity scores and human annotations [35]. Hence, the relative position should also be compared. For all universal text embeddings, the histogram of $Sim\_12$ moves towards the right (more similar) from data in group 1 (dissimilar text pairs) to data in group 5 (similar text pairs). This result indicates that the universal text embeddings based semantic similarity is correlated with the human annotations.
@@
The histogram of $Sim\_1neg1$ within each group is illustrated as orange color in Figure 2: the semantic similarity values are within the range [0.6, 1] for all groups and all universal text embeddings. In order to answer how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings, the histogram of $Sim\_1neg1$ is compared with $Sim\_12$ within in each group. For Group 1: The human annotated similarity between S1 and S2 is within [0, 0.2], which means they are very dissimilar. From the first row of Figure 2, it can be seen that S1 is more similar to its negated version NegS1 (orange color) than to S2 (blue color) across all universal text embeddings (in 99.27% cases on average). Similar observations can be found for Group 2 and 3. For Group 4: The human annotated similarity between S1 and S2 is within [0.6, 0.8] which means they are roughly similar. From the fourth row of Figure 2, it can be seen that the similarity distribution of $Sim\_1neg1$ overlaps with $Sim\_12$. On average (across all embedding models), S1 is more similar to NegS1 than S2 in 51.57% cases. In Group 5: The human annotated similarity between S1 and S2 is within [0.8, 1] which means they are very similar. From the last row of Figure 2, it can be seen that S1 is more similar to S2 than NegS1 (in 77.25% cases). These observations reveal that the distribution of $Sim\_12$ and $Sim\_1neg1$ overlaps most within Group 4, suggesting that these embeddings lack negation awareness, interpreting negated text pairs as roughly similar in terms of semantic meaning. This aligns with the definition of human-annotated semantic similarity scores, as Group 4 has scores ranging from 0.6 to 0.8, indicating that the negated and original sentences are equivalent in terms of topic and syntax, but differ in the presence of negation. This suggests that topic and syntactic information are prioritized over negation information in defining semantic similarity for these embeddings. However, for NLI or sentiment analysis tasks, understanding negation is crucial as it can reverse sentiment entirely. Therefore, adjusting the trade-off among different semantic information such as topic and negation presents a significant challenge for universal text embeddings in achieving task universality.
IV Improve negation awareness in universal text embeddings

### v2

# Semantic Adapter for Universal Text Embeddings: Diagnosing and Mitigating Negation Blindness to Enhance Universality

@@ Abstract
Text embeddings are crucial for various natural language processing tasks, gaining popularity in both industry and academia. Recent progress in training data and LLMs has greatly enhanced the development of universal text embeddings, aiming for a single unified model that can handle diverse tasks, domains and languages. However, due to biases inherent in popular evaluation benchmarks, certain perspectives or abilities of these models remain unassessed. One such overlooked aspect is the models’ capacity for negation awareness. To address this gap in the existing literature, this paper presents a comprehensive analysis of the negation awareness in state-of-the-art universal text embedding models. Our investigation reveals a substantial deficiency in these models, as they frequently interpret negated text pairs as semantically similar. This flaw undermines their effectiveness in accurately understanding negated statements. To mitigate this issue and enhance the universality of text embeddings, we introduce a lightweight, parameter-free negation adapter. The proposed solution involves a data-efficient and computational-efficient embedding re-weighting method that does not require modifications to the parameters of existing text embedding models. The proposed solution is able to improve text embedding models’ negation awareness significantly on both simple negation understanding task and complex negation understanding task. Furthermore, the proposed solution can also significantly improve the negation awareness of Large Language Model based task-specific high dimensional universal text embeddings. These results not only bridge a key “universality gap” but also pave the way for a modular semantic adapter paradigm toward more universal, robust, and environmentally conscious text embeddings.
[…]
123
1 Introduction
@@
Text embedding has gained significant attention from both industry and academia due to its crucial role in various Natural Language Processing (NLP) tasks such as information retrieval [37], sentiment analysis [41, 51], text clustering [48, 38], etc. Recently, the importance of text embeddings has been further highlighted in the context of Large Language Models (LLMs) based applications, such as Retrieval Augmented Generation (RAG) systems. This is primarily because these applications rely on high-quality text embeddings for performing vector search, which involves retrieving the most relevant contexts/documents for LLM Question Answering (QA) tasks [10].
@@
The field of text embeddings has recently achieved significant advancements, transitioning from contextual embeddings like BERT [16] to universal text embeddings. These universal models aim to create a comprehensive embedding framework capable of handling varied input lengths, downstream tasks, domains, and languages. This progress can be attributed to the advances in the quantity and quality of diverse text datasets across different tasks [47, 4], synthetic data generated by different LLMs [26, 45], using LLMs as backbones and the emergence of benchmarks with the focus on novel task and domain generalization such as the Massive Text Embedding Benchmark (MTEB) [33]. Representative universal embedding models such as GTE [29], BGE [47], E5 [44, 45, 46], Gecko [26] or LLM2Vec [5] can effectively address diverse downstream tasks without fine-tuning. They have demonstrated significant improvements over contextual text embeddings across diverse tasks including information retrieval, reranking, clustering and pair classification tasks [10]. However, evaluation benchmarks such as MTEB are biased towards tasks such as retrieval, classification and clustering [33, 10, 11].
Despite the widespread adoption of popular evaluation benchmarks, these frameworks often fail to assess certain critical capabilities of text embedding models. In particular, the negation awareness remains underexamined: a fundamental linguistic construct that plays a pivotal role in understanding sentiment and meaning [3]. Its misrepresentation can alter meaning, distort sentiment analysis, and invalidate inference in applications such as opinion mining and fact verification [3, 39, 21]. Previous works have shown that contextual text embeddings like BERT [16] lack understanding of negations and fail to attribute sufficient importance to the word “not” [18, 20, 21]. This bias can compromise the model’s output, thereby undermining the reliability of these systems, particularly in scenarios involving opinion mining, fact-checking, or assessing answer correctness [3]. However, the extent of this “negation blindness” in universal embeddings has not yet been systematically studied.
@@
  • We formalize negation as a semantic stress test for universal embeddings, constructing controlled paraphrase–negation contrast sets to examine negation awareness across embedding models.
  • We introduce the negation adapter, a data and compute efficient re-weighting module that amplifies negation-sensitive dimensions without any fine-tuning of the embedding models.
@@
2 Related Works
2.1 Universal text embeddings
The field of text embeddings has seen substantial improvements in recent years. Text embedding models can be categorized in four eras [10]: the first era stands for count-based embeddings such as Bag of Words [19] and Term Frequency-Inverse Document Frequency (TF-IDF) [31] which are transformed into low-dimensional dense embeddings with methods like Latent Semantic Indexing (LSA) [15]; the second era is static dense word embeddings represented by Word2Vec [32], GloVe [35] and FastText [6] which can not deal with polysemy; the third era is contextualized embeddings where context-sensitive dynamic embeddings can adapt or change based on context and deal with polysemy, represented by Embeddings From Language Models (ELMo) [34], Generative Pre-trained Transformer (GPT) [36] and Bidirectional Encoder Representations from Transformers (BERT) [16]. Recently, there has been significant progress in the fourth generation of universal text embeddings which aim at building a unified comprehensive text embedding to address a multitude of input text length, tasks, domains and languages [10].
@@ 2.2 Negation understanding
Negation plays an important role in many NLP tasks including Natural Language Inference (NLI) [3], sentiment analysis [39], Natural Language Understanding (NLU) tasks [21], etc. The complexity of negation, as highlighted by [39], presents substantial challenges for commercial sentiment analysis. Contextual text embedding models such as BERT [16] is shown to be insensitive to negations in [18] through a completion task involving masked hypernyms in affirmative and negated versions of each sentence. Other Contextual text embeddings including Transformer-XL [14] and ELMO [34] are also found to exhibit poor performance in distinguishing between affirmative and negative sentences when subjected to negative cloze tests [24]. Moreover, they tend to make identical predictions regardless of the polarity of the sentence [43]. In the study conducted by [20, 21], a novel benchmark was created to evaluate the ability of contextual embedding models in terms of recognizing negation. The experimental findings revealed that even after fine-tuning with additional negation-containing pairs, contextual text embeddings (RoBERTa [30], XLNet [50] and BERT [16]) still face challenges in accurately understanding negation.
Several studies have been conducted in order to improve the negation awareness of contextual text embedding models such as BERT. In [22], BERTNOT is proposed by fine-tuning a pre-trained BERT model using a knowledge distillation objective as well as a novel unlikelihood objective, which is based on negated generic sentences extracted from a raw text corpus. The proposed BERTNOT has been observed to effectively incorporate negation knowledge into the BERT-base model. However, it has been found to be ineffective with the BERT-large model. A novel negation-focused pre-training strategy is proposed in [42] to improve the performance and generalization ability of negation detection using targeted data augmentation and negation masking. Another new pre-training framework with a weighted cross-entropy loss and elastic weight consolidation regularization was proposed in [40] to overcome catastrophic forgetting and augment the negation knowledge in contextual embeddings.
@@
3 Universality Gap Diagnosis
@@
Semantic similarity refers to the degree of overlap in meaning [13]. Measuring the semantic similarity using cosine over text embeddings is one of the most adopted method [33]. Text embedding based semantic similarity is also used as evaluation metrics to measure the answer correctness of RAGs or LLMs [17, 1]. Despite its wide usage, there is few study trying to understand what kind of semantic information is encoded and highlighted in the semantic similarity based on text embeddings. In the literature of Semantic Textual Similarity (STS) studies such as [2, 13], semantic similarity is defined as (re-scaled to the range between 0 and 1):
  • 0 means the pair of texts are on different topics;
  • 0.2 means the pair of texts are not equivalent, but are on the same topic;
  • 0.4 means the pair of texts are not equivalent, but share some details;
@@
  • 0.6 means the pair of texts are roughly equivalent, but some important information differs/missing;
@@
  • 0.8 means the pair of texts are mostly equivalent, but some unimportant details differ;
  • 1 means the pair of texts are completely equivalent;
@@
One limitation of this definition is that: the semantic similarity is mainly based on topic information while other semantic information such as negation is not included. When it comes to two pieces of text where one is the negated version of the other, the situation becomes more complex. For instance, consider these two sentences: (1) "The horse is white." and (2) "The horse is not white.": from a purely syntactic perspective, the sentences are very similar. They share the same subject and predicate, with the only difference being the presence of negation in (2). From a purely topic perspective, the sentences are similar as they share similar topics: the color of the horse. However, semantic textual similarity is not just about the structure or the words used, but about the meaning conveyed by the sentences. The negation in the second sentence changes the meaning entirely. According to [40], the significance of negation in understanding the semantic relationship in natural language texts is profound, as it reverses the overall semantic meaning of a sentence. Consequently, we formalize negation as a stress test of embedding universality and robustness in this section to study how universal text embeddings interpret the semantic similarity between negated text pairs.
3.1 Dataset
@@
STSB [13] is a collection of English datasets that have been used in the SemEval and *SEM STS shared tasks over the period from 2012 to 2017 [13]. Annotation of text pair’s similarity is accomplished via crowdsourcing, integrating pragmatic and global knowledge, thereby enhancing their interpretability and utility for subsequent tasks [13]. Within the scope of this study, the human annotation similarity scores in the STSB dataset, which typically range from 0 to 5, are re-scaled to the range between 0 and 1. Given STSB text pairs S1 ($sentence\_1$) and S2 ($sentence\_2$) with the human annotated semantic similarity in the form of [$S1_{i}$, $S2_{i}$, $Sim_{i}$] (where $i$ is from 1 to N), a sentence-level negation tool (focuses on verbal negations and supports the simple addition and deletion of negation cues) developed by [3] is used to create $NegS1_{i}$: the simple negated version of $S1_{i}$ (e.g. transform "The horse is white" into its simple negated version of "The horse is not white"). The training partition of the dataset is used for analysis (in this section) and training (in next section), while the development and test sets are combined for testing (in next section).
3.2 Universal text embeddings
@@ 3.3 Experiments
In order to study how universal text embeddings interpret negation, the training data of STSB are separated into five groups based on the human annotated similarity $Sim_{i}$ between $S1_{i}$ and $S2_{i}$: group 1 (0-0.2), group 2 (0.2-0.4), group 3 (0.4-0.6), group 4 (0.6-0.8), and group 5 (0.8-1). The semantic similarity $Sim\_12$ between S1 and S2 as well as the semantic similarity $Sim\_1neg1$ between S1 and its negated version NegS1 within each group based on each universal text embeddings are calculated:
$\begin{split}Sim\_12_{i}=Cos(\textbf{e}(S1_{i}),\textbf{e}(S2_{i}))\\ Sim\_1neg1_{i}=Cos(\textbf{e}(S1_{i}),\textbf{e}(NegS1_{i}))\end{split}$ (1)
where $\textbf{e}()$ is the embedding model. By comparing $Sim\_1neg1$ with $Sim\_12$ relatively within each group, we can get the insights on how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings.
@@
The histogram of $Sim\_12$ within each group is illustrated in blue color in Figure 2: row 1 to row 5 represents the results from group 1 to group 5 and each column represents a text embedding model. Generally speaking, the absolute similarity values do not match the human annotation. For instance, the human annotation for group 1 is between 0 and 0.2, while the similarity values calculated by most embedding models are between 0.2 and 0.8. For E5 series, similarity values are between 0.6 and 1 for group 1. This substantial discrepancy highlights that the raw cosine similarity scores generated by these models lack interpretability, as they are not calibrated to scales of human-judged similarity. Consequently, a cosine value of 0.5 cannot be interpreted as "moderately similar" across different models, as each model’s absolute scale is arbitrary. On the other hand, the evaluation metric used by STSB is Pearson correlation between embedding based semantic similarity scores and human annotations [13]. Hence, the relative position should also be compared. For all universal text embeddings, the histogram of $Sim\_12$ moves towards the right (more similar) from data in group 1 (dissimilar text pairs) to data in group 5 (similar text pairs). This result indicates that the universal text embeddings based semantic similarity is correlated with the human annotations.
@@
The histogram of $Sim\_1neg1$ within each group is illustrated as orange color in Figure 2: the semantic similarity values are within the range [0.6, 1] for all groups and all universal text embeddings. In order to answer how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings, the histogram of $Sim\_1neg1$ is compared with $Sim\_12$ within in each group. For Group 1: The human annotated similarity between S1 and S2 is within [0, 0.2], which means they are very dissimilar. From the first row of Figure 2, it can be seen that S1 is more similar to its negated version NegS1 (orange color) than to S2 (blue color) across all universal text embeddings (in 99.27% cases on average). Similar observations can be found for Group 2 and 3. For Group 4: The human annotated similarity between S1 and S2 is within [0.6, 0.8] which means they are roughly similar. From the fourth row of Figure 2, it can be seen that the similarity distribution of $Sim\_1neg1$ overlaps with $Sim\_12$. On average (across all embedding models), S1 is more similar to NegS1 than S2 in 51.57% cases. In Group 5: The human annotated similarity between S1 and S2 is within [0.8, 1] which means they are very similar. From the last row of Figure 2, it can be seen that S1 is more similar to S2 than NegS1 (in 77.25% cases). These results reveal a pronounced universality gap: the distribution of $Sim\_12$ and $Sim\_1neg1$ overlaps most within Group 4, suggesting that these embeddings interpreting negated text pairs as roughly similar in terms of semantic meaning.
4 Negation Adapter: A Lightweight Alternative to Fine-Tuning
Our universality gap diagnosis (Section 3) demonstrates that state-of-the-art universal embeddings exhibit “negation blindness”, underscoring the need for interventions to enhance negation sensitivity for downstream tasks that rely on the semantic understanding of negation. Previous studies on contexual text embeddings [22, 42, 40] assume that these embedding models lack of negation knowledge and focus on improving negation awareness by fine-tuning pre-trained models on negation datasets. However, fine-tuning a universal embedding model on a negation-specific dataset runs counter to the very definition and goals of “universal” text embeddings. In particular:
  • Universality vs Specialization: a universal embedding aims to perform robustly across diverse input lengths, tasks, domains etc. Fine-tuning on negation data may distort representations for other semantic features (e.g. topics, syntax, etc.), undermining its general-purpose utility and risking performance degradation on other axes.
  • Fine-tuning large backbones (often hundreds of millions or billions of parameters) on specialized data is time and energy intensive. This repeated retraining for each new identified universality gap undermines sustainability and eco-friendly deployment. Additionally, each fine-tuned variant necessitates storage, versioning, and integration into pipelines with increasing engineering overhead.
  • High-quality large size negation datasets are rare and expensive to construct for fine-tuning tasks.